OpenStudy (anonymous):

What is the simplified form of (x-5)/(x^2-3x-10) multiplied by (x+2)/(x^2+x-12)? Can anyone solve this problem???

OpenStudy (anonymous):

doesn't really mean multiply it means "factor and cancel"

OpenStudy (anonymous):

could you please guide me through this problem I have tried solving it multiple times and cant seem to solve it.

OpenStudy (anonymous):

did you factor both denominators?

OpenStudy (anonymous):

\[\frac{x-5}{(x-5)(x+2)}\times \frac{x+2}{(x-3)(x+4)}\]

OpenStudy (anonymous):

Yes i got (x-5)(x+2) and (x-3)(x+4)

OpenStudy (anonymous):

Would the answer be (1)/(x-3)(x+4)

OpenStudy (anonymous):